OpenStudy (anonymous):

Show that the function y(t)= 1/3t^2 +c/t is a solution to the different eqution ty' +y=t^2, then find the particular solution that corresponds to the intial condition y(1)=2

OpenStudy (anonymous):

@Hero

OpenStudy (irishboy123):

you can plug the solution into the DE so calculate \(t(\frac{1}{3}t^2 +\frac{c}{t})' + (\frac{1}{3}t^2 +\frac{c}{t})\) , or you can solve this directly using, say, an integrating factor.

OpenStudy (anonymous):

sorry I still do not understand it.

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

starting with \(y = \frac{1}{3}t^2 +\frac{c}{t}\) calculate \(y' = ...\) then evaluate ty' + y [alternatively note that \((ty)' = ty' + y\) ... but i suspect you should ignore that thought and just go as advised]

OpenStudy (anonymous):

oay so y'= 2/3t^2 -c/t^2

OpenStudy (anonymous):

what will be my ty' in this case?? @IrishBoy123

OpenStudy (freckles):

well your y' is just a little off

OpenStudy (freckles):

\[y=\frac{1}{3}t^{2}+\frac{c}{t}=\frac{1}{3}t^2+ct^{-1} \\ y'=\frac{1}{3} \cdot (2t)+c \cdot (-1 t^{-2}) =\frac{2}{3}t-\frac{c}{t^2}\]

OpenStudy (freckles):

anyways to find ty' you just multiply t to the y' there

OpenStudy (freckles):

\[t y'=t(\frac{2}{3}t-\frac{c}{t^2}) \text{ you can distribute }\]

OpenStudy (freckles):

and then of course you add that that to your y you have and see if the result is t^2

OpenStudy (anonymous):

ty'= 2/3 t^2 - c/t^3 + y where Y is 1/3t^2 +c/t 4/3t^2 -c/t^3+c/t

OpenStudy (freckles):

t/t^2 is 1/t not 1/t^3

OpenStudy (anonymous):

oh i see so it would be +c/t -c/t so that term is cancled

OpenStudy (anonymous):

and am left with 4/3t^2

OpenStudy (freckles):

\[y=\frac{1}{3}t^2+\frac{c}{t} \\ y'=\frac{2}{3}t-\frac{c}{t^2} \\ ty'=\frac{2}{3}t^2-\frac{c}{t} \\ \text{ so } ty'+y=\frac{2}{3}t^2-\frac{c}{t}+\frac{1}{3}t^2+\frac{c}{t}\]

OpenStudy (freckles):

well 2+1 is 3 not 4

OpenStudy (freckles):

so it should be 3/3 t^2

OpenStudy (anonymous):

sorry. okkay so I am left with 3/3t^2 which is t^2

OpenStudy (freckles):

right which means that thing that they called y is like totally a solution to the differential equation given

OpenStudy (freckles):

since both sides of the diff equation are equal when we plug that in

OpenStudy (anonymous):

right but at the end the questin has another part which is then find the particular soulution.... y(1)=2

OpenStudy (freckles):

then you can find c by using the condition given

OpenStudy (anonymous):

so I plug in 1 where t is in the original equation?

OpenStudy (freckles):

yep and after that replace y(1) since you replace all the t's with 1 ...replace y(1) with 2 and solve equation for c

OpenStudy (freckles):

\[y(1)=\frac{1}{3}(1)^2+\frac{c}{1}\] replaced t's with 1 now replace y(1) with 2 since y(1)=2

OpenStudy (anonymous):

2=1/3 +c 2-1/3=c 6-1/3=c 5/3=c

OpenStudy (freckles):

that's right so the particular solution they are looking for is y(t)=1/3 t^2+5/(3t)

OpenStudy (freckles):

I just plugged that value for c back into y(t)=1/3t^2+c/t

OpenStudy (anonymous):

right so my final answer would be y(t)=1/3 t^2+5/(3t)

OpenStudy (freckles):

yep for the particular solution part of the question

OpenStudy (anonymous):

alright yaya done with this! tysm u are really good and thaks for taking the tiem to type u everyting in equation format I love it when ppl at openstudy teach me with that. TYSM :)

OpenStudy (freckles):

well some of it I was lazy on and didn't type in the thingy

OpenStudy (anonymous):

and that fac tthat It takes a while to type up everything in equation format. NO, u did good for most of the part

OpenStudy (freckles):

k thanks glad you understand the problem though :)

OpenStudy (anonymous):

I understand it if it's less sentence and more math. but thank you for helping me!